DeWitt, Corp just paid a dividend of $1 and the dividend will be growing at a constant rate of 25% for 2 years, and after that it will be growing at 5%. What is the intrinsic value of DeWitt's stock if investors require a rate of return of 11.0%? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

Intrinsic Value = ?
Answer #1

D1=(1*1.25)=1.25

D2=(1.25*1.25)=1.5625

Value after year 2=(D2*Growth rate)/(Required rate-Growth rate)

=(1.5625*1.05)/(0.11-0.05)

=27.34375

Hence intrinsic value=Future dividend and value*Present value of discounting factor(rate%,time period)

=1.25/1.11+1.5625/1.11^2+27.34375/1.11^2

=$24.59(Approx).